A 12.4 L cylinder contains 785.5g of ethane at 22.5c. determine the amount of the gas (in gram) that must be released to reduce the pressure to 19.2 atm, assuming it is an ideal gas.
1
Expert's answer
2020-08-17T13:33:22-0400

We need to use ideal gas law here:

PV=nRT

Lets find the initial chemical amount of ethane:

n=m/M=785.5/30=26.1833 mol

The final chemical amount of ethane we will calculate through ideal gas law:

n=PV/RT

n= 9.8191 mol

Thus, we need to remove 26.1833 - 9.8191=16.3642 mol of ethane. In grams it will be 490.9 g.

Answer: 490.9 g